40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 12 22:51:26 2016, Koji, Update, General, Clipping could not be improved by the PZTs - NEED ANOTHER VENT 
    Reply  Thu Oct 13 00:23:54 2016, ericq, Update, VAC, Pre-vent prep 
       Reply  Thu Oct 13 09:30:38 2016, Steve, Update, General, vent 79 is happening vent79precond.pngventing.png
          Reply  Thu Oct 13 16:48:34 2016, Steve, Update, VAC, vent 79 is completed vent_79.png
             Reply  Thu Oct 13 19:28:14 2016, gautam, Update, General, in-air alignment 
                Reply  Fri Oct 14 10:31:53 2016, gautam, Update, General, doors are off ITMY and BS/PRM chambers 
                   Reply  Fri Oct 14 18:33:55 2016, gautam, Update, General, AS clipping investigations 7x
                      Reply  Mon Oct 17 22:45:16 2016, gautam, Update, General, AS beam centered on all OMs 7x
                         Reply  Tue Oct 18 18:56:57 2016, gautam, Update, General, OM5 rotated to bypass OMC, green scatter is from window to PSL table OMCchamber.pdfgreenGhosts.JPGIMG_3322.JPGIMG_3326.JPG
                            Reply  Wed Oct 19 16:41:55 2016, gautam, Update, General, Heavy doors back on 8x
             Reply  Thu Oct 20 09:17:08 2016, Steve, Update, VAC, Pumpdown 80 has started readyPd80.pngbeforePd80.pngpd80at1h.png7d_atm.png
                Reply  Thu Oct 20 15:59:52 2016, Steve, Update, VAC, Pumpdown 80 has completed 7hPd80.png8hPd80.png9hPd80.png
                   Reply  Fri Oct 21 02:06:20 2016, gautam, Update, General, IFO recovery 
                      Reply  Mon Oct 24 11:39:13 2016, gautam, Update, General, ALS recovered ALSOutOfLoop20161024.pdfXendPDHOLTF20161024.pdf
                   Reply  Fri Oct 21 09:28:21 2016, Steve, Update, VAC, Vac Normal reached VacNormal.png
                      Reply  Wed Nov 23 11:32:13 2016, Steve, Update, VAC, TP3 drypump replaced again 
                         Reply  Fri Nov 10 08:52:21 2017, Steve, Update, VAC, TP3 drypump replaced 
                            Reply  Fri Feb 16 15:57:58 2018, Steve, Update, VAC, TP3 drypump replaced 
                Reply  Thu Jul 12 16:18:22 2018, Steve, Update, VAC, Vent preparations before_Vent.pngbefore_Vent_cond.png
                   Reply  Thu Jul 12 23:59:14 2018, gautam, Update, General, Vent objectives and prep 
                      Reply  Fri Jul 13 02:52:11 2018, gautam, Update, General, Vent objectives and prep preVentStatus.png
                         Reply  Fri Jul 13 17:13:45 2018, Koji, Update, General, Vent objectives and prep P_20180713_020603.jpg
                         Reply  Fri Jul 13 18:01:13 2018, gautam, Update, General, Low power MC 
Message ID: 12576     Entry time: Fri Oct 21 02:06:20 2016     In reply to: 12575     Reply to this: 12578
Author: gautam 
Type: Update 
Category: General 
Subject: IFO recovery 

The pressure on the newly installed gauge on the X arm was 6E-5 torr when I came in today evening, so I decided to start the recovery process.

  1. I first tried working at low power. I was able to lock the IMC as well as the arms. But the dither alignment didn't work so well. So I decided to go to nominal PSL power.
  2. I first changed the 2" HR mirror that is used to send all the MC REFL light to the MC REFL PD in low power operation with a 10% BS. I then roughly aligned the beam onto the PD using the tiny steering mirror. At this point, I also re-installed the ND filters on the end Transmon QPDs and also the CCD at the Y end.
  3. I then rotated the waveplate (the second one from the PSL aperture) until I maximized the power as measured just before the PSL shutter with a power meter. I then re-aligned the PMC to maximize transmission. After both these steps, we currently have 1.09W of IR light going into the IMC
  4. I then re-aligned MC REFL onto the PD (~90mW of light comes through to the PD) and maximized the DC output using an oscilloscope. I then reverted the Autolocker to the nominal version from the low power variant that has been running on megatron during the vent (although we never really used it). The autolocker worked well and I was able to lock the IMC without much trouble. I tweaked the alignment sliders for the IMC optics, but wasn't able to improve the transmission much. It is ~14600 cts right now, which is normal I think
  5. I then centered the beams onto the WFS QPDs, ran the WFSoffsets script after turning the inputs to the WFS servos off, and ran the relief script as well - I didn't try anything further with the IMC
  6. I then tried to lock the arms - I first used the green to align the test-masses. Once I was able to lock to a green 00-mode, I saw strong IR flashes and so I was able to lock the Y arm. I then ran the dither. Next, I did the same for the X arm. Even though I ran LSCoffsets before beginning work tonight, the Y arm transmission after maximization is ~5, and that for the X arm is ~2.5. I refrained from running the normalization scripts in case I am missing something here, but the mode itself is clearly visible on the cameras and is a 00-mode.
    GV edit 21Oct2016: For the Y-arm, the discrepancy was down to TRY being derived from the high gain PD as opposed to the QPD. Switching these and running the dither, TRY now maxes out at around 1.0. For TRX, the problem was that I did not install one of the ND filters - so the total ND was 1.2 rather than 1.6, which is what we were operating at and which is the ND on TRY. Both arms now have transmission ~1 after maximizing with the dither alignment...
  7. The AS spot looks nice and round on the camera, although the real check would be to do the sort of scan Yutaro and Koji did, and monitor the ASDC levels. I am leaving this task for tomorrow, along with checking the recycling cavities.
  8. Lastly, I centered the Oplevs for all the TMs

 

ELOG V3.1.3-